20 bags of cement , of 40 kg , cost Rs1800. Find the cost of 15 bags of cement weighing 3 kg each.

Grade:6

1 Answers

Arun
6 years ago
Dear Sindhu
 
20 bafs 40 kg each
Total 20*40 = 800 kg
Thus 800 kg for 1800 rs
 
Now for 1 kg = 1800/800 = 2.25 rs
 
Now 
 
15 bags 3 kg each 
Total weight = 15*3 = 45 kg
 
Hence total cost
= 45 * 2.25 = 101.25 Rs
 
 
Regards
Arun (askIITians forum expert)